ProbabilityHard
Question
Let E and F be two independent events. The probability that both E and F happen is 1/ 12 and the probability that neither E. nor F happen is 1 /2. Then,
Options
A.P(E) = 1/ 3, P(F) = 1/ 4
B.P(E) = 1/ 2, P(F) = 1/ 6
C.P(E) = 1/ 6, P(F) = 1/ 2
D.P(E) = 1/ 4, P(F) = 1/ 3
Solution
Both E and F happen ⇒ P(E ∩ F) =
and neither E nor F happens ⇒ P(
∩
) =
But for independent events, we have
P(E ∩ F) = P(E) P(F) =
......(i)
and P(
∩
) = P(
) P(
)
= {1- P(E)} {1- P(F)}
= 1- P(E) - P(F) + P(E)P(F)
⇒
= 1 - {P(E) + P(F)} + 
⇒ P(E) + P(F) = 1 -
......(ii)
On solving Eqs. (i) and (ii) , we get
either P(E)=
and P(F) =
or P(E) =
and P(F) =
and neither E nor F happens ⇒ P(
∩
) =
But for independent events, we have
P(E ∩ F) = P(E) P(F) =
......(i) and P(
∩
) = P(
) P(
)= {1- P(E)} {1- P(F)}
= 1- P(E) - P(F) + P(E)P(F)
⇒
= 1 - {P(E) + P(F)} + 
⇒ P(E) + P(F) = 1 -
......(ii) On solving Eqs. (i) and (ii) , we get
either P(E)=
and P(F) =
or P(E) =
and P(F) =
Create a free account to view solution
View Solution FreeMore Probability Questions
The probability that an event A happens in one trial of an experiment is 0.4. Three independent trails of the experiment...If two cards are drawn from a pack of card one by one. If first drawn card is replaced then the probability of getting t...A dice is thrown then the odds against of getting the digit 6 is-...The probability that two persons have same date of birth is (in non-leap year)...Two integers $x$ and $y$ are chosen from the set $\{ 0,1,2,3,\ldots,2n - 1,2n\}$, with replacement. The probability that...